OpenStudy (anonymous):

Does anyone know how to simplify complex fractions or simplify the differences?

OpenStudy (asnaseer):

It would be better for you to post your actual question. most people here will be able to help you on almost any topic in maths.

OpenStudy (anonymous):

well it says simplify the difference: (x)over(3x+9) - (8)over(xsquared+3x)

OpenStudy (asnaseer):

so its:\[\frac{x}{3x+9}-\frac{8}{x^2+3x}\]

OpenStudy (anonymous):

yep. How did you do that?

OpenStudy (asnaseer):

first thing you need to do is to factorise the denominators. (you can use the equation editor to enter latex expressions here)

OpenStudy (anonymous):

Oh..okay. thanks. and im not quite sure how to factor 3x+9

OpenStudy (asnaseer):

can you notice anything common between the terms: 3x and 9

OpenStudy (asnaseer):

what divides into both these terms?

OpenStudy (anonymous):

3

OpenStudy (asnaseer):

correct, therefore we can factorise it as follows:\[3x+9=3(x+3)\]try and do the same with the second denominator

OpenStudy (anonymous):

oh. hah okay

OpenStudy (asnaseer):

BTW: if you don't know how to enter latex yet then you can enter your equation: (x)over(3x+9) - (8)over(xsquared+3x) as: x/(3x+9) - 8/(x^2+3x) to make it easier for others to read (and for you to enter)

OpenStudy (anonymous):

so would the second one be x(x+3) and thanks.

OpenStudy (asnaseer):

perfect!

OpenStudy (asnaseer):

so now we can rewrite the equation as follows:\[\frac{x}{3x+9}-\frac{8}{x^2+3x}=\frac{x}{3(x+3)}-\frac{8}{x(x+3)}\]

OpenStudy (anonymous):

then would I multiply the opposite side by the factored denominator?

OpenStudy (asnaseer):

from here you now need to find the common denominator for both fractions which is the least common multiple (LCM) of both.

OpenStudy (anonymous):

its both of them right? since there isnt more than one thats the same.

OpenStudy (asnaseer):

the two denominators are: 3(x+3) and x(x+3) which term is common in both?

OpenStudy (anonymous):

3 and x

OpenStudy (asnaseer):

the 3 and x are the ones that are not common :)

OpenStudy (asnaseer):

(x+3) occurs in both denominators, therefore (x+3) is the common term, therefore you need to include this just once

OpenStudy (asnaseer):

next you multiply it by all the other "uncommon" terms - which are 3 and x

OpenStudy (asnaseer):

so the LCM for the denominator would be 3x(x+3)

OpenStudy (asnaseer):

can you see how both denominators would divide exactly into this LCM?

OpenStudy (anonymous):

so you multiply each side with the lcm?

OpenStudy (anonymous):

No, i'm confused on where i go from writing down (x)/3(x+3) - (8)/x(x+3)

OpenStudy (asnaseer):

it might help to do an example with just numbers. suppose we had:\[\frac{5}{3*4}+\frac{3}{2*4}\]

OpenStudy (asnaseer):

the LCM of the denominators here would be 3*2*4

OpenStudy (asnaseer):

so we would write:\[\frac{5}{3*4}=\frac{5*2}{3*4*2}\]and:\[\frac{3}{2*4}=\frac{3*3}{2*4*3}\]

OpenStudy (asnaseer):

this leads to both fractions having the "same" denominator, so we can then do the following:\[\frac{5*2}{3*4*2}=\frac{10}{24}\]\[\frac{3*3}{2*4*3}=\frac{9}{24}\]\[\frac{5}{3*4}+\frac{3}{2*4}=\frac{10}{24}+\frac{9}{24}=\frac{19}{24}\]

OpenStudy (anonymous):

okay. So i multiply each side with the lcm?

OpenStudy (asnaseer):

that is why we first need to find a "common" denominator

OpenStudy (asnaseer):

ok, so lets summarise where we got to:\[\frac{x}{3x+9}-\frac{8}{x^2+3x}=\frac{x}{3(x+3)}-\frac{8}{x(x+3)}\]and we know LCM is 3x(x+3), therefore:\[\frac{x}{3x+9}-\frac{8}{x^2+3x}=\frac{x}{3(x+3)}-\frac{8}{x(x+3)}\]\[\qquad=\frac{x*x}{3x(x+3)}-\frac{8*3}{3x(x+3)}\]\[\qquad=\frac{x^2}{3x(x+3)}-\frac{24}{3x(x+3)}\]

OpenStudy (asnaseer):

now we have a common denominator so we can get:\[\qquad=\frac{x^2-24}{3x(x+3)}\]

OpenStudy (anonymous):

where did you get \[x ^{2}-24\]?

OpenStudy (asnaseer):

from the working out above where we got to:\[\qquad=\frac{x^2}{3x(x+3)}-\frac{24}{3x(x+3)}\]

OpenStudy (asnaseer):

\[x*x=x^2\]\[8*3=24\]

OpenStudy (anonymous):

so you multiplied each side by 3x(x+3)?

OpenStudy (asnaseer):

not quite, I multiplied each fraction with something that would turn its denominator into the LCM. writing it like this might help:\[\frac{x}{3(x+3)}=\frac{x}{3(x+3)}*\frac{x}{x}=\frac{x^2}{3x(x+3)}\]

OpenStudy (asnaseer):

we can do this because multiplying any number by 1 doesn't change its value, i.e.\[\frac{x}{x}=1\]

OpenStudy (asnaseer):

similarly, we multiplied the second fraction by:\[\frac{3}{3}\]

OpenStudy (anonymous):

sorry i had to get off for a minute.

OpenStudy (asnaseer):

did you follow the explanation above?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok - good :)